Q2. Find the domain and range of each function.

a) 𝑓(𝑑)=π‘₯21+π‘₯2

b) 𝑔(𝑧)=|π‘₯βˆ’3|π‘₯βˆ’3

c) 𝑓(π‘₯)=1√4βˆ’π‘₯2


1
Expert's answer
2021-05-07T11:41:32-0400

In part c, "a^2\geq \frac{1}{4}." Hence the range will beΒ "(\frac{1}{2},\infty)."

In part a), domain is the entire real line since "x^{21}+x^2" is defined over the entire real line.